# Mitigating Suboptimality of Deterministic Policy Gradients in Complex Q-functions

### v1

@@ Appendix A Proof of Convergence of Maximizer Actor in Tabular Settings
To prove convergence, we will show that the sequence of policies ${\mu_{n}}$ yields monotonically non-decreasing value functions that converge to a stable value function $V^{N}$.
Policy Evaluation Converges
@@
Given the current policy $\mu_{n}$, the policy evaluation computes the action-value function $Q^{\mu_{n}}$, satisfying:
$Q^{\mu_{n}}(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s,a,s^{\prime})V^{\mu_{n}}(s^{\prime}),$
where $V^{\mu_{n}}(s^{\prime})=Q^{\mu_{n}}(s^{\prime},\mu_{n}(s^{\prime}))$.
In the tabular setting, the Bellman operator $\mathcal{T}^{\mu_{n}}$ defined by
$[\mathcal{T}^{\mu_{n}}Q](s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s,a,s^{\prime})Q(s^{\prime},\mu_{n}(s^{\prime}))$
is a contraction mapping with respect to the max norm $\|\cdot\|_{\infty}$ with contraction factor $\gamma$,
$\|\mathcal{T}^{\mu_{n}}Q-\mathcal{T}^{\mu_{n}}Q^{\prime}\|_{\infty}\leq\gamma\|Q-Q^{\prime}\|_{\infty}.$
Therefore, iteratively applying $\mathcal{T}^{\mu_{n}}$ converges to the unique fixed point $Q^{\mu_{n}}$.
Policy Improvement with DPG and Maximizer Actor
Step 1: DPG Update
We define $\tilde{\mu}_{n}$ as the DPG policy that locally updates $\mu_{n}$ towards maximizing the expected return based on $Q^{\mu_{n}}$. For each state $s$, we perform a gradient ascent step using the Deep Policy Gradient (DPG) method to obtain an improved policy $\tilde{\mu}_{n}$:
$\tilde{\mu}_{n}(s)\leftarrow\mu_{n}(s)+\alpha\nabla_{a}Q^{\mu_{n}}(s,a)\big{|}_{a=\mu_{n}(s)},$
@@
$V^{\tilde{\mu}_{n}}(s)\geq V^{\mu_{n}}(s),\quad\forall s\in\mathcal{S}.$
@@
Since the sequence $\{V^{\mu_{n}}\}$ is monotonically non-decreasing and bounded above by $V^{*}$, it converges to some $V^{\infty}\leq V^{*}$. Given the finite number of possible policies, the sequence $\{\mu_{n}\}$ must eventually repeat a policy. Suppose that at some iteration $N$, the policy repeats, i.e., $\mu_{N+1}=\mu_{N}$.
At this point, since the policy hasn’t changed, we have:
$\mu_{N}(s)=\arg\max_{a\in\{\tilde{\mu}_{N}(s),\nu_{1}(s),\dots,\nu_{k}(s)\}}Q^{\mu_{N}}(s,a),\quad\forall s\in\mathcal{S}.$
Since $\tilde{\mu}_{N}(s)$ is obtained by performing a DPG update on $\mu_{N}(s)$, and we have that $\mu_{N}(s)$ maximizes $Q^{\mu_{N}}(s,a)$ among $\{\tilde{\mu}_{N}(s),\nu_{1}(s),\dots,\nu_{k}(s)\}$, it must be that:
$Q^{\mu_{N}}(s,\mu_{N}(s))\geq Q^{\mu_{N}}(s,a),\quad\forall a\in\{\tilde{\mu}_{N}(s),\nu_{1}(s),\dots,\nu_{k}(s)\}.$
Moreover, since $\tilde{\mu}_{N}(s)$ is obtained via gradient ascent from $\mu_{N}(s)$, and yet does not yield a higher $Q$-value, it implies that:
$\nabla_{a}Q^{\mu_{N}}(s,a)\big{|}_{a=\mu_{N}(s)}=0.$
This suggests that $\mu_{N}(s)$ is a local maximum of $Q^{\mu_{N}}(s,a)$. This shows that this modification to the policy iteration algorithm of DPG is guaranteed to converge.
Since the set $\{\tilde{\mu}_{N}(s),\nu_{1}(s),\dots,\nu_{k}(s)\}$ includes more actions from $\mathcal{A}$, $\mu_{N}(s)$ is the action that better maximizes $Q^{\mu_{N}}(s,a)$ than $\tilde{\mu}_{N}$. Therefore, $\mu_{N}$ is a greedier policy with respect to $Q^{\mu_{N}}$ than $\tilde{\mu}_{N}$.
Appendix B Proof of Reducing Number of Local Optima in Successive Surrogates
Theorem B.1.
Consider a state $s\in\mathcal{S}$, the function $Q$ as defined in Eq. 1, and the surrogate functions $\Psi_{i}$ as defined in Eq. 7. Let $N_{\text{opt}}(f)$ denote the number of local optima (assumed countable) of a function $f:\mathcal{A}\to\mathbb{R}$, where $\mathcal{A}$ is the action space. Then,
$N_{\text{opt}}(Q(s,a))\geq N_{\text{opt}}(\Psi_{0}(s,a;\{a_{0}\}))\geq N_{\text{opt}}(\Psi_{1}(s,a;\{a_{0},a_{1}\}))\geq\dots\geq N_{\text{opt}}(\Psi_{k}(s,a;\{a_{0},\dots,a_{k}\})).$
@@ Appendix B Proof of Reducing Number of Local Optima in Successive Surrogates
For each $i\geq 0$, define the surrogate function $\Psi_{i}$ recursively:
$\Psi_{i}(s,a;\{a_{0},\dots,a_{i}\})=\max\left\{Q(s,a),\tau_{i}\right\},$ (11)
where
$\tau_{i}=\max_{0\leq j\leq i}Q(s,a_{j}).$
Note that $\tau_{i}$ is non-decreasing with respect to $i$, i.e., $\tau_{i+1}\geq\tau_{i}$.
Our goal is to show that for each $i\geq 0$,
$N_{\text{opt}}(\Psi_{i}(s,a;\{a_{0},\dots,a_{i}\}))\geq N_{\text{opt}}(\Psi_{i+1}(s,a;\{a_{0},\dots,a_{i+1}\})).$
We proceed by considering how the set of local optima changes from $\Psi_{i}$ to $\Psi_{i+1}$.
@@
Consider any local optimum $a^{\prime}$ of $\Psi_{i}$. There are two cases:
Case 1: $Q(s,a^{\prime})>\tau_{i}$
In this case, $\Psi_{i}(s,a^{\prime})=Q(s,a^{\prime})$ and $\Psi_{i}$ coincides with $Q$ in a neighborhood of $a^{\prime}$. Since $a^{\prime}$ is a local optimum of $\Psi_{i}$, it is also a local optimum of $Q$. Because $\tau_{i+1}\geq\tau_{i}$, there are two subcases:
Subcase 1a: $Q(s,a^{\prime})>\tau_{i+1}$
Here, $\Psi_{i+1}(s,a^{\prime})=Q(s,a^{\prime})$ and, in a neighborhood of $a^{\prime}$, $\Psi_{i+1}$ coincides with $Q$. Thus, $a^{\prime}$ remains a local optimum of $\Psi_{i+1}$.
Subcase 1b: $Q(s,a^{\prime})\leq\tau_{i+1}$
Since $Q(s,a^{\prime})>\tau_{i}$ and $\tau_{i+1}\geq\tau_{i}$, this implies $\tau_{i+1}>\tau_{i}$ and $Q(s,a^{\prime})=\tau_{i+1}$. Then,
$\Psi_{i+1}(s,a^{\prime})=\tau_{i+1},$
and in a neighborhood around $a^{\prime}$, $\Psi_{i+1}(s,a)\geq\tau_{i+1}$. Thus, $a^{\prime}$ is not a local optimum of $\Psi_{i+1}$ because there is no neighborhood where $\Psi_{i+1}(s,a)<\Psi_{i+1}(s,a^{\prime})$.
Case 2: $Q(s,a^{\prime})\leq\tau_{i}$
In this case, $\Psi_{i}(s,a^{\prime})=\tau_{i}$, and $\Psi_{i}$ is constant at $\tau_{i}$ in a neighborhood of $a^{\prime}$. Thus, $a^{\prime}$ may be considered a local optimum in $\Psi_{i}$ if the function does not exceed $\tau_{i}$ nearby. When moving to $\Psi_{i+1}$, since $\tau_{i+1}\geq\tau_{i}$, we have:
$\Psi_{i+1}(s,a^{\prime})=\tau_{i+1}\geq\tau_{i}.$
In the neighborhood of $a^{\prime}$, $\Psi_{i+1}$ remains at least $\tau_{i+1}$, so $a^{\prime}$ is not a local optimum in $\Psi_{i+1}$ unless $Q(s,a)<\tau_{i+1}$ in a neighborhood around $a^{\prime}$.
However, since $\Psi_{i+1}(s,a)\geq\tau_{i+1}$ for all $a$, the function does not decrease below $\Psi_{i+1}(s,a^{\prime})$ in any neighborhood of $a^{\prime}$. Therefore, $a^{\prime}$ is not a local optimum of $\Psi_{i+1}$.
Conclusion: From the above cases, we observe that:
  • Any local optimum $a^{\prime}$ of $\Psi_{i}$ where $Q(s,a^{\prime})>\tau_{i+1}$ remains a local optimum in $\Psi_{i+1}$.
  • Any local optimum $a^{\prime}$ of $\Psi_{i}$ where $Q(s,a^{\prime})\leq\tau_{i+1}$ does not remain a local optimum in $\Psi_{i+1}$.
Since $\Psi_{i+1}$ does not introduce new local optima (because $\Psi_{i+1}(s,a)\geq\Psi_{i}(s,a)$ for all $a$ and coincides with $Q$ only where $Q(s,a)>\tau_{i+1}$), the number of local optima does not increase from $\Psi_{i}$ to $\Psi_{i+1}$.
Base Case: For $i=0$, we have:
$\Psi_{0}(s,a;\{a_{0}\})=\max\left\{Q(s,a),Q(s,a_{0})\right\}.$
If we consider $Q(s,a_{0})$ to be less than the minimum value of $Q(s,a)$ (which can be arranged by choosing $a_{0}$ appropriately or by defining $\tau_{0}$ to be less than $\inf_{a}Q(s,a)$), then $\Psi_{0}(s,a)=Q(s,a)$, and the base case holds trivially.
Inductive Step: Assuming that
$N_{\text{opt}}(\Psi_{i}(s,a;\{a_{0},\dots,a_{i}\}))\leq N_{\text{opt}}(\Psi_{i}(s,a;\{a_{0},\dots,a_{i}\})),$
we have shown that
$N_{\text{opt}}(\Psi_{i+1}(s,a;\{a_{0},\dots,a_{i+1}\}))\leq N_{\text{opt}}(\Psi_{i}(s,a;\{a_{0},\dots,a_{i}\})).$
By induction, it follows that:
$N_{\text{opt}}(Q(s,a))\geq N_{\text{opt}}(\Psi_{0}(s,a;\{a_{0}\}))\geq N_{\text{opt}}(\Psi_{1}(s,a;\{a_{0},a_{1}\}))\geq\dots\geq N_{\text{opt}}(\Psi_{k}(s,a;\{a_{0},\dots,a_{k}\})).$
Appendix C Environment Details
Figure 13: Benchmark Environments involve discrete action space tasks like Mine World and recommender systems (simulated and MovieLens-Data) and restricted locomotion tasks.
Figure 14: Mining Expedition. The red agent must reach the green goal by navigating the grid and using one or more pick-axes to clear each mine blocking the path.
C.1 MiningEnv
@@
The grid world environment, introduced in  §5, requires an agent to reach a goal by navigating a 2D maze as soon as possible while breaking the mines blocking the way.
@@
    Agent Position: A 2D vector representing the agent’s x and y coordinates.
@@
    Agent Direction: One dimension representing directions (0: right, 1: down, 2: left, 3: up).
@@
    Surrounding Path: A 4-dimensional vector indicating if the adjacent cells are empty or a goal (1: empty/goal, 0: otherwise).
@@
    Front Cell Type: A $(K+1)$-dimensional one-hot vector with first $K$ dimensions representing the type of mine and the last dimension representing if the cell is empty (zero) or goal (one).
Finally, we will normalize each dimension to [0, 1] with each dimension’s minimum/maximum value.
Termination: An episode terminates when the agent reaches the goal or after 100 timesteps. Upon reset, the grid layout changes while keeping the agent’s start and goal positions fixed.
Actions: Actions include navigation (up, down, left, right) and pick-axe categories. Navigation changes the agent’s direction and attempts to move forward. The agent cannot step into a mine but will change direction when trying to step onto a mine or the border of the grid. The pick-axe tool actions (50 types) have a predefined one-to-one mapping of how they interact with the mines, which means they can be successfully applied to only one kind of mine, and either transform that kind of mine into another type of mine or directly break it.
Reward: The agent’s reward comprises a goal-reaching reward, a distance-based step reward, and rewards for successful tool use or mine-breaking. The goal reward is discounted by steps taken over the episode, encouraging shorter paths to reach the goal.
@@
Action Representations Actions are represented as a 4D vector with normalized values [0, 1] as described in Figure 14. Dimensions represent skill category (navigation or pick-axe), movement direction (right, down, left, up), mine type where this action can be applied, and the outcome of applying the tool to the mine, respectively.
C.2 RecSim
In the simulated recommendation system (RecSys) environment, the agent selects an item from a large set that aligns with the user’s interests. Users are modeled with dynamically changing preferences that evolve based on their interactions (clicks). The agent’s objective is to infer these evolving preferences from user clicks and recommend the most relevant items to maximize the total number of clicks.
State: The user’s interest is represented by an embedding vector $e_{u}\in\mathbb{R}^{n}$, where $n$ is the number of item categories. This embedding evolves over time as the user interacts with different items. When a user clicks on an item with embedding $e_{i}\in\mathbb{R}^{n}$, the user interest embedding $e_{u}$ is updated as follows:
$e_{u}$ $\leftarrow e_{u}+\Delta e_{u},\quad\text{with probability }\frac{e_{u}^{\top}e_{i}+1}{2}$
$e_{u}$ $\leftarrow e_{u}-\Delta e_{u},\quad\text{with probability }\frac{1-e_{u}^{\top}e_{i}}{2},$
where $\Delta e_{u}$ represents an adjustment that depends on the alignment between $e_{u}$ and $e_{i}$. This update mechanism adjusts the user’s preference towards the clicked item, reinforcing the connection between the current action on future recommendations.
Action: The action set consists of all items that can be recommended, and the agent must select the item most relevant to the user’s long-term preferences over the episode.
Reward: The reward is based on user feedback: either a click (reward = 1) or skip (reward = 0). The user model computes a score for each item using the dot product of the user and item embeddings:

### v2

@@ 9 Proof of Convergence of Maximizer Actor in Tabular Settings
9.1 Policy Evaluation Converges
Given a deterministic policy $\pi$ (in our case $\pi=\mu_{M}$), the policy evaluation computes the action-value function $Q^{\pi}$, which satisfies the Bellman equation:
$Q^{\pi}(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s,a,s^{\prime})Q^{\pi}(s^{\prime},\pi(s^{\prime})).$
In the tabular setting, the Bellman operator $\mathcal{T}^{\pi}$ defined by
$\mathcal{T}^{\pi}Q](s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s,a,s^{\prime})Q(s^{\prime},\pi(s^{\prime}))$
is a contraction mapping with respect to the max norm $\|\cdot\|_{\infty}$ with contraction factor $\gamma$:
$\|\mathcal{T}^{\pi}Q-\mathcal{T}^{\pi}Q^{\prime}\|_{\infty}\leq\gamma\|Q-Q^{\prime}\|_{\infty}.$
@@ 9.1 Policy Evaluation Converges
9.2 Policy Improvement with DPG and Maximizer Actor
At iteration $n$, suppose we have a policy $\mu_{n}$.
Step 1: Policy Evaluation
Compute $Q^{\mu_{n}}$ by solving:
$Q^{\mu_{n}}(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s,a,s^{\prime})Q^{\mu_{n}}(s^{\prime},\mu_{n}(s^{\prime})).$
Step 2: Policy Improvement
(a) DPG Update
Perform a gradient ascent step using the Deep Policy Gradient (DPG) method to obtain an improved policy $\tilde{\mu}_{k+1}$:
$\tilde{\mu}_{k+1}(s)=\mu_{n}(s)+\alpha\nabla_{a}Q^{\mu_{n}}(s,a)\big|_{a=\mu_{n}(s)},$
@@ 9.2 Policy Improvement with DPG and Maximizer Actor
$V^{\tilde{\mu}_{k+1}}(s)\geq V^{\mu_{n}}(s),\quad\forall s\in\mathcal{S}.$
@@
Since $\{V^{\mu_{n}}\}$ is bounded above by $V^{*}$ (the optimal value function), it converges. In a finite MDP, there are only finitely many possible policies. Thus, the sequence $\{\mu_{n}\}$ must eventually repeat, and because each policy improvement is non-decreasing, the policies stabilize at an optimal policy $\mu^{*}$.
Theorem 4.1 demonstrates how the maximizer actor at least improves over the DPG policy. Yet, there is no guarantee that this achieves the global optimum because of the implicit dependence on the policy gradient algorithm. Therefore, Theorem 4.1 is limited to showing convergence in the presence of a maximizer actor to a local optimum and shows that our algorithm is a stable RL algorithm — not one that is globally optimal. It also shows how the locally optimal policy learned with a maximizer actor might improve the locally optimal policy learned by DPG.
10 Proof of Reducing Number of Local Optima in Successive Surrogates
Theorem 10.1.
Consider a state $s\in\mathcal{S}$, $Q$ in Eq. 1, and $\Psi_{i}$ in Eq. 7. Let $N_{\text{opt}}(f)$ be the number of local optima (assumed countable) of a function $f:\mathcal{A}\to\mathbb{R}$, where $\mathcal{A}$ is the action space. Then,
$N_{\text{opt}}(Q(s,a))\geq N_{\text{opt}}(\Psi_{0}(s,a;\{a_{0}\})),\dots,\geq N_{\text{opt}}(\Psi_{k}(s,a;\{a_{0},\dots,a_{k}\}))$
@@ 10 Proof of Reducing Number of Local Optima in Successive Surrogates
Consider two consecutive surrogate functions $\Psi_{i}(s,a;\{a_{0},\dots,a_{i}\})$ and $\Psi_{i+1}(s,a;\{a_{0},\dots,a_{i+1}\})$,
$\Psi_{i}(s,a;a_{<i})=\max\left\{Q(s,a),\max_{j<i}Q(s,a_{j})\right\},$
$\Psi_{i+1}(s,a;a_{<i+1})=\max\left\{Q(s,a),\max_{j<i+1}Q(s,a_{j})\right\},$
Let $\tau_{i}=\max_{j<i}Q(s,a_{j})$ and $\tau_{i+1}=\max_{j<i+1}Q(s,a_{j})$.
@@
Consider a given state $s$ and any particular local optimum in $\Psi_{i}$ at $a^{\prime}$, there can be two cases:
  1. 1.
    If Q(s, a’) > $\tau_{i+1}$, then $\Psi_{i+1}(s,a^{\prime};a_{<i+1})=Q(s,a^{\prime})$.
    Since, a’ is a local optimum of $\Psi_{i}$, there exists $\epsilon>0$ $\Psi_{i}(s,a^{\prime}\pm\epsilon;a_{<i})=Q(s,a^{\prime}\pm\epsilon)<\Psi_{i}(s,a^{\prime};a_{<i})=Q(s,a^{\prime})$
    Therefore, $\Psi_{i+1}(s,a^{\prime}\pm\epsilon;a_{<i+1})=Q(s,a^{\prime}\pm\epsilon)<\Psi_{i+1}(s,a^{\prime};a_{<i+1})=Q(s,a^{\prime})$ Thus, a’ is also a local optimum of $\Psi_{i+1}$.
  2. 2.
    If $Q(s,a^{\prime})\leq\tau_{i+1}$, then $\Psi_{i+1}(s,a^{\prime};a_{<i+1})=\tau_{i+1}$, and there exists $\epsilon>0$, such that $\Psi_{i+1}(s,a^{\prime}\pm\epsilon;a_{<i+1})=\tau_{i+1}$. Thus, a’ is not a local optimum of $\Psi_{i+1}$
Finally, $\Psi_{i+1}$ does not add any new local optima, because $\tau_{i+1}\geq\tau_{i}$ and thus all points where $\Psi_{i+1}(s,a;a_{<i+1})=Q(s,a)$, we have $\Psi_{i}(s,a;a_{<i})=Q(s,a)$. Therefore $\forall i\geq 1$,
$N_{\text{opt}}(\Psi_{i}(s,a;\{a_{0},\dots,a_{i}\}))\geq N_{\text{opt}}(\Psi_{i+1}(s,a;\{a_{0},\dots,a_{i+1}\})$
The same analysis extends for $Q$ and $\Psi_{1}$, by substituting $\tau_{0}<\min{Q}$ to be a very small value. Thus, by induction, we have,
$N_{\text{opt}}(Q(s,a))\geq N_{\text{opt}}(\Psi_{0}(s,a;\{a_{0}\})),\dots,\geq N_{\text{opt}}(\Psi_{k}(s,a;\{a_{0},\dots,a_{k}\}))$
11 Environment Details
Figure 12: This figure provides the visual description of the environment setup.
11.1 MiningEnv
@@
The grid world environment, introduced in Sec. 5, requires an agent to reach a goal by navigating a 2D maze as soon as possible while breaking the mines blocking the way.
@@
    Agent Position: Agent Position occupies two dimensional of the vector. The first dimension represents the x-axis value, and the second one represents the y.
@@
    Agent Direction: It only takes one channel with value [0, 1, 2, 3]. Each number represents one direction, and they are 0-right, 1-down, 2-left, and 3-up.
@@
    Surrounding Path: This information takes four channels. Each represents whether the cell in that direction is an empty cell or a goal.
@@
    Front Cell Type: This information is in one-hot form and occupies the last K + 1-dimensional vector, which provides the information of which kind of mine is in front of the agent. If the front cell is an empty cell or the goal, the $(K+1)^{th}$ channel will be one, and others remain to be zero
Ultimately, we will normalize each dimension to [0, 1] with each dimension’s minimum/maximum value. Each time we reset the environment, the layout of the whole grid world will be changed, except for the agent start position and the goal position.
Termination: An episode is terminated in success when the agent reaches the goal or after a total of 100 timesteps.
Actions: The base action set combines two kinds of actions: navigation actions and pick-axe(tool) actions. The navigation action set is a fixed set, which contains four independent actions: going up, down, left, and right, corresponding with the direction of the agent. They will change the agent’s direction first and then try to make the agent take one step forward. Note that, different from the empty cell, the agent cannot step onto the mine, which means that if the agent is trying to take a step towards a mine or the border of the world, then the agent will stay in the same location while the direction will still be changed. Otherwise, the agent can step onto that cell. An agent will succeed if it reaches the goal position. The size of the pick-axe action set is equal to 50. Each tool has a one-to-one mapping, which means they can and only can be successfully applied to one kind of mine, and either transform that kind of mine into another type of mine or directly break it.
Reward: The agent receives a large goal reward for reaching the goal. The goal reward is discounted based on the number of action steps taken to reach that location, thus rewarding shorter paths. To further encourage the agent to reach the goal, a small exploration reward is added whenever the agent gets closer to the goal, and a negative equal penalty is added whenever the agent gets further to the goal. And also, when the agent successfully applies a tool, it will gain a small reward. When the agent successfully breaks a mine, it will also gain a small bonus.
@@
Action Representations: The action representations are 4-dimensional vectors manually defined using a mix of number ids, and each dim is scaled to [0, 1]. as shown in Graph 13. Dimensions 1 identifies the category of skills (navigation, pick-axe), 2 distinguishes movement skills (right, down, left, up), 3 denotes the mine on which this tool can be successfully applied, and 4 shows the result of applying this tool. We will normalize the action embedding space to [0, 1] for each dimension.
Figure 13: Mining Env Setting Description
11.2 RecSim
The simulated RecSys environment requires an agent to select an item that match the user’s interest out of a large item-set. We simulate users with a dynamically changing preference upon clicks. Thus, the agent’s task is to infer this preference from user clicks and recommend the most relevant item to maximize a total number of clicks.
State: The user interest embedding ($e_{u}\in\mathbb{R}^{n}$ where $n$ denotes the number of categories of items) represents the user interest in categories that transitions over time as the user consumes different items upon click. So, when the user clicks an item with the corresponding item embedding($e_{i}\in\mathbb{R}^{n}$; the same $n$ as the one for the user embedding) then the user interest embedding($e_{u}$) will be updated as follows;
$\Delta(e_{u})$ $=(-y|e_{u}|+y)\cdot(1-e_{u}),\text{ for }y\in[0,1]$
$e_{i}$ $\leftarrow e_{u}+\Delta(e_{u})\text{ with probability}[e_{u}^{T}e_{i}+1]/2$
$e_{u}$ $\leftarrow e_{u}-\Delta(e_{u})\text{ with probability}[1-e_{u}^{T}e_{i}]/2$
This essentially pulls the user’s preference towards the item that was clicked.
Action: The action set contains many recommendable items. So, the agent has to find the most relevant item to a user given the item-set. See below regarding how these items are represented.
Reward: The base reward is a simulated user feedback (e.g., clicks). The user model (Ie et al., 2019) stochastically skips or clicks the recommended item based on the present user interest embedding ($e_{u}$). Concretely, the user model computes the following score on the recommended item;